OpenStudy (anonymous):

a body starts sliding over horizontal surface with an initial velocity of .5 m/s. due to friction its velocity decreases .05 m/s. how mch time will it take to stop

OpenStudy (anonymous):

10 s

OpenStudy (anonymous):

10seconds

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

v = u + at 0.5 = 0.05t t=10s

OpenStudy (anonymous):

Details plzz

OpenStudy (anonymous):

v =0 becoz the body will stop finally

OpenStudy (anonymous):

so 0 = u - at since it is deccelerating -u=-at

OpenStudy (anonymous):

a = 0.05

OpenStudy (anonymous):

@Rohangrr u got it

OpenStudy (anonymous):

lol u'll explain and i'll not get

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

nthing